OpenStudy (anonymous):

Solve. 2x^2 + x = 15 please, algebra 2 Solve. 2x^2 + x = 15 please, algebra 2 @Mathematics

OpenStudy (anonymous):

2x^2+x-15 = 0

OpenStudy (anonymous):

2x^2 + x = 15 2x^2 + x - 15=0, now solve for x using quadratic formula or by fActoring

jhonyy9 (jhonyy9):

this is equal (x+3)(2x-5)

jhonyy9 (jhonyy9):

x+3=0 x=-3 and 2x-5=0 2x=5 x=5/2

jhonyy9 (jhonyy9):

ok ?